OpenStudy (anonymous):

The scale on a map is 1 inch: 500 miles. If two cities are 875 miles apart, how far are they on the map? help please.

OpenStudy (bibby):

Set up a proportion or use dimensional analysis. \(\huge \frac{1in}{500}=\frac{x in}{875}\) 875=500x solve for x \(\huge \frac{1in}{500\cancel{mi}}*\frac{875\cancel{mi}}{}\)

OpenStudy (anonymous):

Thank you!